# Quantifying and Mitigating Unimodal Biases in Multimodal Large Language Models: A Causal Perspective

### v3

@@ 7 Conclusion
This paper presents a comprehensive approach to quantifying and mitigating the unimodal biases in MLLMs. Through our causal inference framework, we provide an in-depth analysis to assess the causal effects of such biases on the model’s prediction in VQA problems. The introduced MORE dataset challenges MLLMs to engage in multi-hop reasoning and to overcome language and vision biases, thereby pushing the boundaries of their reasoning capabilities. Our proposed solutions, including the DeVA framework and the fine-tuned LLaVA models, demonstrate significant potential in enhancing the reasoning abilities of MLLMs.
8 Limitations
Our current generation of rationales is based on heuristic rules. Experiments in Section 5.3 have demonstrated the effectiveness of incorporating rationales into instructions. Therefore, we believe that refining and polishing these rationales with an LLM (e.g., ChatGPT) could be beneficial. Besides, the Wikidata5M dataset we employed was released in 2021, and some information in the knowledge graph may be outdated. Although we have made efforts to manually verify the test set and try to ensure it does not contain incorrect information, it is still inevitable that errors may occur within the extensive training data.
@@ A.1 Question Generation
We present the prompt template for generating language bias options in Section 3.3 in Figure 7.
@@ A.2 Language Bias Option Generation
We present the prompt template for generating language bias options in Section 3.3 in Figure 8.
A.3 Training Instance with Causal Rationale
Figure 9: Prompt template of the training instance in MORE.
Figure 10: Human evaluation results of MORE.
An example of the training instance in MORE is shown in Figure 9.
@@ C.2 Human Evaluation
We conduct a human evaluation of 100 questions randomly chosen from the MORE dataset to validate and assess the quality of the generated questions. This evaluation is carried out by three human annotators, who are provided with detailed guidelines and illustrative examples before starting the evaluation process. For each question, given the visual context and ground truth answer, the annotators are asked to determine whether: 1) the question is valid, 2) the question has a possible answer alternative, 3) the question does not match the answer, or 4) the question is unclear or confusing. For each sample, we ask three human annotators to judge each generated question into one of the four options.
The outcomes of this human evaluation are summarized in Figure 10. The results are encouraging, with 90.6% of the generated questions being classified as valid by the annotators, further demonstrating the quality of our datasets.
Appendix D Baselines and Datasets Details
D.1 Baselines
@@
1) BLIP2 Li et al. (2023), which utilizes a scalable multimodal pre-training method to enable LLMs to understand images. We employ its BLIP2-OPT Zhang et al. (2022)-6.7B variant.
@@
2) InstructBLIP Dai et al. (2023), an instruction-tuned version of BLIP-2 on various tasks including VQA. We employ its InstructBLIP-Vicuna Chiang et al. (2023)-13B variant.
@@
3) mPLUG-Owl Ye et al. (2023), which proposes a new two-stage training method for aligning images and text. We employ its mPLUG-Owl-Llama Touvron et al. (2023a)-7B variant.
@@
4) LLaVA Liu et al. (2023a), which translates images into texts of captions and bounding boxes, and prompts GPT-4 to generate a multimodal instruct-tuning dataset in the context of seed examples. We employ its LLaVA-Llama Touvron et al. (2023a)-13B variant.
D.2 Datasets
OK-VQA Marino et al. (2019) is a VQA dataset that necessitates external knowledge beyond the visible content in images. It is crafted by crowdsourced participants who form complex questions based on extensive information available on Wikipedia.
INFOSEEK Chen et al. (2023) is another VQA dataset aimed at information-seeking questions that require more than just commonsense knowledge. It combines human-annotated questions with visual entity recognition datasets and Wikidata to generate complex question-answer pairs.
Appendix E Implementation Details of Intervention
In this section, we introduce the implementation details of interventions on questions and images. Given the impracticality of enumerating all possible perturbations of $T$ and $R$, practical results within a specific subset of $T$ and $R$ could be attained by intervening on $Q$ while leaving $S$ unaffected, and intervening on $I$ while leaving $E$ unaffected, respectively Stolfo et al. (2022).
For questions, in the analysis of $\operatorname{TCE}$, the image stays the same while the question is altered to a different one from the dataset that relates to the same entity, ensuring a ground truth shift post-intervention. In the analysis of $\operatorname{DCE}$, we similarly keep the image constant and let ChatGPT rephrase the question, altering its textual form but preserving its semantic meaning (verified through manual evaluation), thereby maintaining an unchanged ground truth after the intervention.
For images, in the analysis of $\operatorname{TCE}$, we maintain a constant question while replacing the image with another from the dataset that corresponds to the same question but features different entities, ensuring a change in the ground truth post-intervention. In the analysis of $\operatorname{DCE}$, the question remains unchanged, but the image is replaced with one of the same entity from an alternative source, specifically using Google Lens’s API***Web interface available at https://images.google.com.) to keep the ground truth consistent after the intervention.
Appendix F The DeVA Framework
@@
In this section, we first present the overall framework, then we will go over each part of it in detail.
F.1 Overall Framework
@@
Given a question $Q$ and an image $I$, the VQA task demands the system to return an output $A$ that concisely answers the question. As shown in Figure 11, we first initialize a question decomposer $D$ to analyze $Q$ and break it down into manageable subquestions. Then, we employ a verifier $V$ to confirm the accuracy of the original answer to each decomposed subquestion. The generation verifier typically involves active information-seeking and answer-verification, which acquires the necessary context or information needed to investigate and revise the answers. This includes two optional operations: image retrieval to seek images similar to $I$ and determine their titles, or text retrieval with a specific query to fetch pertinent documents and summarize their content. This iterative process of answering and verifying will continue until we resolve each subquestion. Finally, the verified answer $A$ is output following the aforementioned reasoning process and retrieved information.
F.2 Question Decomposer
@@
For a given VQA problem, MLLMs often simply exploit a spurious path to make predictions due to the inherent language and vision biases. In order to alleviate this issue, motivated by Chain-of-Thought reasoning Wei et al. (2022), we encourage MLLMs to decompose the question $Q$ before outputting the answer, so as to gradually solve a complex question that requires multi-hop reasoning. As shown in Figure 11, for the question “Which country hosted the next World Cup after this venue?”, our decomposer breaks it down into two subquestions :
@@
Such decomposition will explicitly constrain the model to comprehend and extract the truth semantics of the question, thus avoiding simply exploring a spurious path to give the answer.
F.3 Verifier
@@
Although our framework is applicable to any image retrieval method, in this paper, we mainly utilize Google Image Search to obtain a broad range of information related to the image as provided by Google Lens API. This information encompasses various details, such as knowledge graph entities and captions of analogous or identical images. The availability of these details can vary based on the image input provided to Google Image Search. Then, the verifier gleans relevant information from captions associated with visually similar images, so as to verify the original answer and conduct the next round of reasoning.
@@
Figure 11: An overview of our proposed Decompose-Verify-Answer (DeVA) Framework.
Figure 12: Prompt template of DeVA framework.
F.4 Final Verified Answer
Finally, the improved response that takes verification into account is generated. This is executed by a final prompt where the context takes into account all of the previous reasoning steps, the baseline response, and the verification question-answer pairs, so that the corrections can take place. The prompt template is shown in Figure 12.
Appendix G Implementation Details of Fine-tuning LLaVA
We adopt LLaVA-13B-v1.5 as the base model and employ the LoRA Hu et al. (2022) technique. During pre-training, only LoRA parameters are optimized. We set the rank of LoRA modules to 128. Our model is optimized with a learning rate of 2e-4 and a linear warm-up for the first 3% steps. The batch size is 16 and the number of epochs is 1. All the LoRA parameters are trained on 8 NVIDIA A100 GPUs with 80GB memory. We fine-tune LLaVA on the training sets of INFOSEEK and MORE, respectively. For MORE, we incorporate the generated rationale into the instructions of some samples, so as to enhance the models’ reasoning ability. The template of a training instance can be found in Figure 9.
Appendix H Case Study
We conduct a case study on the development set of MORE in Figure 13$\sim$16, including both the “Open-ended” and “Multi-Choice” settings.
@@ Appendix H Case Study
Figure 14: Two-hop case (multi-choice setting) in the dev set of MORE. The wrong answers are marked in red and the correct is in green.

### v4

@@ 7 Conclusion
This paper presents a comprehensive approach to quantifying and mitigating the unimodal biases in MLLMs. Through our causal inference framework, we provide an in-depth analysis to assess the causal effects of such biases on the model’s prediction in VQA problems. The introduced MORE dataset challenges MLLMs to engage in multi-hop reasoning and to overcome language and vision biases, thereby pushing the boundaries of their reasoning capabilities. Our proposed CAVE method demonstrates significant potential in enhancing the reasoning abilities of MLLMs.
Limitations
Our current generation of rationales is based on heuristic rules. Previous works have demonstrated the effectiveness of incorporating rationales into instructions Wei et al. (2022). Therefore, we believe that refining and polishing these rationales with an LLM (e.g., ChatGPT) could be beneficial. Besides, the Wikidata5M dataset we employed was released in 2021, and some information in the knowledge graph may be outdated. Although we have made efforts to manually verify the test set and try to ensure it does not contain incorrect information, it is still inevitable that errors may occur within the extensive training data.
Acknowledgments
We thank all the anonymous reviewers for their valuable feedback throughout the review process. This work is supported in part by Ucap Cloud and the State Key Laboratory of General Artificial Intelligence.
@@ A.1 Question Generation
We present the prompt template for generating language bias options of Section 3.2 in Figure 7.
@@ A.2 Language Bias Option Generation
We present the prompt template for generating language bias options of Section 3.2 in Figure 8.
@@ C.2 Human Evaluation
Questions
We conduct a human evaluation of 100 questions randomly chosen from the MORE dataset to validate and assess the quality of the generated questions. This evaluation is carried out independently by three human annotators, who are provided with detailed guidelines and illustrative examples before starting the evaluation process. For each question, given the visual context and ground truth answer, we first ask two junior annotators to determine whether: 1) the question is valid, 2) the question allows for an alternative answer, 3) the question does not match the answer, or 4) the question is unclear or ambiguous. If the choices of the two annotators are inconsistent, a senior annotator checks the answers and makes the final decision. The average inter-annotator agreement is 88.6% (Cohen’s kappa).
Rationales
We also conduct a human evaluation of the causal rationales, following the same procedure as described above. The difference is that here we provide annotators with only two options: to assess whether the generated rationales are valid.
As shown in Figure 9, the results are encouraging, with 91% questions and 98% rationales being classified as valid by the annotators, further demonstrating the quality of our datasets.
Appendix D Baselines
@@
1) InstructBLIP Dai et al. (2023), an instruction-tuned version of BLIP-2 on various tasks including VQA. We employ its InstructBLIP-Vicuna Chiang et al. (2023)-13B variant.
@@
2) mPLUG-Owl Ye et al. (2023), which proposes a new two-stage training method for aligning images and text. We employ its mPLUG-Owl-Llama Touvron et al. (2023a)-7B variant.
@@
3) LLaVA Liu et al. (2023a), which translates images into texts of captions and bounding boxes, and prompts GPT-4 to generate a multimodal instruct-tuning dataset in the context of seed examples. We employ its LLaVA-Llama Touvron et al. (2023a)-13B variant.
4) Qwen-VL Bai et al. (2023b), which builds upon Qwen Bai et al. (2023a) and employ multi-stage training pipeline. Qwen-VL facilitates grounding and text comprehension by aligning image-caption-box tuples. It processes inputs of images, text, and bounding boxes, and produces corresponding text and bounding boxes as outputs.
Appendix E Details of the CAVE Framework
@@
In this section, we first present the overall framework as shown in Figure 10, and then we will go over each part of it in detail.
E.1 Overall Framework
@@
Given a question $Q$ and an image $I$, the VQA task demands the system to return an output $A$ that concisely answers the question. As shown in Figure 10, we first initialize a question decomposer $D$ to analyze $Q$ and break it down into manageable subquestions. For each subquestion, we introduce a causality-enhanced reasoner $C$ to evaluate the correctness of the decomposed subquestion. Then, we employ a verifier $V$ to confirm the accuracy of the original answer to each subquestion. The generation verifier typically involves active information-seeking and answer-verification, which acquires the necessary context or information needed to investigate and revise the answers. This includes two optional operations: image retrieval to seek images similar to $I$ and determine their titles, or text retrieval with a specific query to fetch pertinent documents and summarize their content. This iterative process of answering and verifying will continue until we resolve each subquestion. Finally, the verified answer $A$ is output following the aforementioned reasoning process and retrieved information.
Figure 10: An overview of our proposed CAVE Framework.
E.2 Question Decomposer
@@
For a given VQA problem, MLLMs may simply exploit a spurious shortcut to make predictions. In order to alleviate this issue, motivated by Chain-of-Thought reasoning Wei et al. (2022), we encourage MLLMs to decompose the question $Q$ before outputting the answer, so as to gradually solve a complex question that requires multi-hop reasoning. As shown in Figure 10, for the question “Which country hosted the next World Cup after this venue?”, our decomposer breaks it down into two subquestions :
@@
Such decomposition will explicitly constrain the model from comprehending and extracting the truth semantics of the question, thus avoiding simply exploring a spurious path to give the answer.
E.3 Causality-enhanced Reasoner
The accuracy of sub-question decomposition plays a crucial role in the subsequent reasoning process. To address this, we introduce causality-enhanced self-reflection, which explicitly guides the model in validating the sub-question it has decomposed. Specifically, based on the proposed causal framework, we guide the model to assess whether the decomposed subquestion changes under two conditions: 1) rephrasing the question text, or 2) replacing the image with another captured by the same entity from a different angle or at a different time. If the model identifies that the decomposed subquestion needs to be altered, this indicates that its initial understanding is influenced by irrelevant factors or biases, failing to capture the true semantics. In such cases, we will enforce a re-decomposition of the question.
E.4 Verifier
@@
Although our framework is applicable to any image retrieval method, in this paper, we mainly utilize Google Image Search to obtain a broad range of information related to the image as provided by Google Lens API [^1] Web interface available at https://images.google.com.. This information encompasses various details, such as knowledge graph entities and captions of analogous or identical images. The availability of these details can vary based on the image input provided to Google Image Search. Then, the verifier gleans relevant information from captions associated with visually similar images, so as to verify the original answer and conduct the next round of reasoning.
@@
E.5 Final Verified Answer
Finally, the improved response that takes verification into account is generated. This is executed by a final prompt where the context takes into account all of the previous reasoning steps, the baseline response, and the verification question-answer pairs, so that the corrections can take place.
Appendix F Case Study
We conduct a case study on the development set of MORE in Figure 11$\sim$14, including both the “Open-ended” and “Multi-Choice” settings.
@@ Appendix F Case Study
Figure 12: Two-hop case (multi-choice setting) in the dev set of MORE. The wrong answers are marked in red.